# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v7

@@ 3.3 Information Processing and Learning Optimization in Models
The fundamental objective of model learning is to more accurately process user input information $Iq$ and to deliver high-quality output information $Ir$ that meets user requirements. Since the modeling and realization of input information $Iq$ are almost identical to those of $It$, as discussed in detail in Section 3.1, they will not be reiterated here. Thus, our focus shifts to the process whereby the updated information $Iu$ within model M processes $Iq$, outputs $Ir$, and is further optimized through learning to yield $Iv$. In this context, we continue to define $Iq$, $Ir$, and $Iv$ in the manner of equations (3.1-3.3).
Definition 10 (Formal Representation of State Processing): Let the states $Sz(z,T_{z})$, $Su(u,T_{u})$, and $Sv(v,T_{v})$ be interpretations $E_{z}$, $E_{u}$, and $E_{v}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, corresponding to object $z$ over time interval $T_{z}$, object $u$ over $T_{u}$, and object $v$ over $T_{v}$, all satisfying Postulate 5. We introduce the following specific functions and predicates in $\mathcal{L}$:
  • Function: $f_{process}^{2}(X,Y)$ denotes the set of well-formed formulas output by state set $X$ after processing state set $Y$. Similarly, when both $X$ and $Y$ satisfy the logical consistency requirement of Postulate 5, $f_{process}^{2}(X,Y)$ also meets this logical consistency requirement.
  • Predicate: $\text{IsSubsetOf}^{2}(X,Y)$ indicates that set $X$ is a subset of set $Y$; $\text{Processable}^{2}(X,Y)$ indicates that state set $X$ is capable of processing state set $Y$ and generating an output; $\text{IsFormulaOf}^{2}(x,X)$ denotes that $x$ is a well-formed formula within state set $X$; and $\text{Eq}^{2}(x,y)$ indicates that $x$ is equal to $y$.
Then suppose $\varphi z\in Sz(z,T_{z})$, $\sup(T_{z},T_{u})\leq\inf(T_{v})$, and
$\varphi z=($ $\forall\Phi z)(\forall\Phi u)(\text{IsSubsetOf}^{2}(\Phi z,Sz(z,T_{z}))\land\text{IsSubsetOf}^{2}(\Phi u,Su(u,T_{u}))\land$
$\text{Processable}^{2}(\Phi z,\Phi u))\rightarrow\exists\varphi v(\text{IsFormulaOf}^{2}(\varphi v,Sv(v,T_{v}))\land$
$\text{Eq}^{2}(\varphi v,f_{process}^{2}(\Phi z,\Phi u)))$ (3.13)
we say that $Sz(z,T_{z})$ is capable of processing $Su(u,T_{u})$ and outputting $Sv(v,T_{v})$; equivalently, we say that the interpretation $E_{z}$ has the ability to process $E_{u}$ and generate $E_{v}$ as output.
Clearly, by combining Definitions 8 and 10, we can also formally express the process whereby state $S_{ou}(ou,Tu_{h})$ processes and learns from $S_{oq}(oq,Tq_{h})$, outputs $S_{or}(or,Tr_{h})$, and is subsequently optimized into $S_{ov}(ov,Tv_{h})$. In addition, when $S_{ou}(ou,Tu_{h})$ and $S_{oq}(oq,Tq_{h})$ both satisfy Postulate 5, the logical consistency of $f_{learn}^{2}(X,Y)$ and $f_{process}^{2}(X,Y)$ ensures that $S_{or}(or,Tr_{h})$ and $S_{ov}(ov,Tv_{h})$ also satisfy Postulate 5. Thus, we have:
Definition 11 (Implementation of Model Information Processing and Learning Optimization): For an AI model M with updated information $Iu$, input information $Iq$, output information $Ir$, and optimized information $Iv$, if all four are recoverable, and the ontological state $S_{ou}(ou,Tu_{h})$ is capable of processing and learning from $S_{oq}(oq,Tq_{h})$, outputting $S_{or}(or,Tr_{h})$, and further optimizing into $S_{ov}(ov,Tv_{h})$, while the carrier state $S_{M}(M,Tu_{h})$ is similarly capable of processing and learning from $S_{cq}(cq,Tq_{h})$, outputting $S_{cr}(cr,Tr_{h})$, and further optimizing into $S_{M}(M,Tv_{h})$, then M is said to possess the capability to realize the information processing and learning optimization of $Iu$ with respect to $Iq$, outputting $Ir$ and optimizing into $Iv$. In short, M possesses information processing and learning optimization capability.
For the neural network model $Mn$ discussed in the previous section, the processable predicate $\text{Processable}^{2}$ is defined as:
$\text{Processable}^{2}(\Phi u,\Phi q)=\text{InputMatch}(\Phi u,\Phi q)\land\text{ActivationFeasible}(\Phi u,\Phi q)\land\text{OutputDefined}(\Phi u)$ (3.14)
[…]
The processing function $f_{process}^{2}$ is defined as:
$f_{process}^{2}(\Phi u,\Phi q)=f_{\Phi u}(\Phi q)$ (3.15)
where $f_{\Phi u}$ is the prediction function explicitly expressed by the structure and parameters of the neural network $Mn$.
Thus, for the neural network model $Mn$, the ontological state of the updated information $Iu$ is $S_{ou}(ou,Tu_{h})=\Phi u$, and the ontological state of the input information $Iq$ is $S_{oq}(oq,Tq_{h})=\Phi q$. When the processable condition $\text{Processable}^{2}(\Phi u,\Phi q)$ is satisfied, the ontological state of the output information $Ir$ is
$S_{or}(or,Tr_{h})=\Phi r=f_{process}^{2}(\Phi u,\Phi q)=f_{\Phi u}(\Phi q)$ (3.16)
demonstrating that the ontological state $S_{ou}(ou,Tu_{h})$ possesses the ability to process $S_{oq}(oq,Tq_{h})$ and output $S_{or}(or,Tr_{h})$. Similarly, since $Iu$, $Iq$ and $Ir$ are all recoverable information, when their ontological states are mapped onto $Mn$ and its interface, the carrier states $S_{Mn}(Mn,Tu_{h})$ also possess the ability to process $S_{cq}(cq,Tq_{h})$ and output $S_{cr}(cr,Tr_{h})$. Thus, $Mn$ is capable of realizing the processing of $Iu$ and $Iq$ to produce $Ir$. Furthermore, according to section 3.2, $Mn$ is also capable of learning from $Iq$ and optimizing into $Iv$.
Naturally, one can just as easily use the formal framework of Definitions 8 and 10 to demonstrate that other typical models—such as linear regression and decision trees—also possess both information processing and learning optimization capability. It is thus evident that this axiomatic and formalized architecture transcends specific algorithms or models, providing a more universal, unified, and interpretable foundational meta-framework for machine learning theory.
@@ 4 Several Important Theorems and Proofs
The Meta-Framework for Machine Learning Theory (MLT-MF) applies to a series of pivotal issues in the field of artificial intelligence, providing many unique and practically valuable theoretical principles and methodological guidelines.
@@ 4.1 Theorem of Model Interpretability
Interpretability of complex models, such as those in deep learning, has long been one of the key challenges in artificial intelligence. However, the academic community has yet to establish a rigorous mathematical formulation of model interpretability [6, 30, 31, 32]. In fact, the interpretability of any phenomenon in the universe is inherently relative. For example, Kepler’s three laws explain the motion of planets around the sun, but such explanations are meaningless to those without elementary mathematical knowledge. Even for the same individual, the explanations that can be understood and accepted may evolve as their knowledge expands. Therefore, interpretability should be understood as a concept that depends on the domain of discourse and the corresponding logical system. Furthermore, it should also be considered in relation to the specific inputs and outputs of the model. Based on MLT-MF, we define the following concepts:
@@
Definition 12 (Model interpretability): Let the state of a model M over time interval $Tm$ be denoted as $S_{M}(M,Tm)$. Given an input state $S_{cq}(cq,Tq_{h})$ and an output state $S_{cr}(cr,Tr_{h})$, suppose there exists a formal system $\mathcal{L}$ and three corresponding interpretations $E_{M}$, $E_{q}$, and $E_{r}$ over finite domains $D_{M}$, $D_{q}$, and $D_{r}$, respectively, such that:
  • $E_{M}$ has the capacity to process $E_{q}$ and produce $E_{r}$ as output, and
  • $S_{M}(M,Tm)$, $S_{cq}(cq,Tq_{h})$, and $S_{cr}(cr,Tr_{h})$ are each bijective enabling mappings of $E_{M}$, $E_{q}$, and $E_{r}$, respectively.
@@
Then, we say that model M is interpretable in $\mathcal{L}$ with respect to the output $S_{cr}(cr,Tr_{h})$ under the input condition $S_{cq}(cq,Tq_{h})$.
In this way, we define model interpretability via the formal system $\mathcal{L}$ and its interpretation. This definition aligns both with the intuitive understanding of "interpretation" and with the theoretical need to characterize the internal logic of models through a series of propositions, making it well-suited for the formal analysis and in-depth study of AI interpretability.
In fact, according to Theorem 1, if a model M has the ability to process an input state $S_{cq}(cq,Tq_{h})$ and produce an output state $S_{cr}(cr,Tr_{h})$, then by Definition 10, the states $S_{M}(M,Tm)$, $S_{cq}(cq,Tq_{h})$, and $S_{cr}(cr,Tr_{h})$ are interpretations $E_{M}$, $E_{q}$, and $E_{r}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, all satisfying Postulate 5. In this case, the mappings $E_{M}\rightrightarrows S_{M}(M,Tm)$, $E_{q}\rightrightarrows S_{cq}(cq,Tq_{h})$, and $E_{r}\rightrightarrows S_{cr}(cr,Tr_{h})$ are self-mappings, and therefore naturally satisfy the condition of being bijective enabling mappings. This leads to:
Corollary 1 (Intrinsic Model Interpretability): If a model M possesses the ability to process an input state $S_{cq}(cq,Tq_{h})$ and produce an output state $S_{cr}(cr,Tr_{h})$, then there exists a formal system $\mathcal{L}_{M}$ in which M is interpretable with respect to the output $S_{cr}(cr,Tr_{h})$ under the input condition $S_{cq}(cq,Tq_{h})$. Since $\mathcal{L}_{M}$ serves as the formal system that interprets model M, it is also referred to as the intrinsic formal system of the model.
As Hegel famously asserted, "What is real is rational." The intrinsic interpretability of a model means that, insofar as it can process inputs and generate outputs according to established rules and procedures, it necessarily possesses an internal logic that is, in principle, interpretable. However, this logic may be highly implicit and abstract, making it difficult for people—even domain experts—to identify and understand. This highlights the need to map and translate this internal logic into a more intuitive and universal logical system, thereby enabling deeper decomposition and analysis of the causal mechanisms, and achieving broader understanding and acceptance. This is the essence of model interpretability, and is precisely the core idea behind Definition 12: to interpret the internal logic of models within more explicit formal systems.
Theorem 3 (Model interpretability and information recoverability): Let model M be capable of processing the information $Iq=\langle oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq}\rangle$ and producing output information $Ir=\langle or,Tr_{h},S_{or},cr,Tr_{m},S_{cr}\rangle$, through its operation described by $Iu=\langle ou,Tu_{h},S_{ou},M,T_{m},S_{M}\rangle$. If all three pieces of information $Iu$, $Iq$, and $Ir$ are recoverable, and their ontological states are interpretations of well-formed formulas in a common formal system $\mathcal{L}$, then model M is interpretable in $\mathcal{L}$ with respect to the output $S_{cr}(cr,Tr_{h})$ under the input condition $S_{cq}(cq,Tq_{h})$.
Proof: According to Definition 5, since $Iu$, $Iq$, and $Ir$ are recoverable, the following are all bijective enabling mappings: $S_{ou}(ou,T_{h})\rightrightarrows S_{M}(M,T_{m})$, $S_{oq}(oq,Tq_{h})\rightrightarrows S_{cq}(cq,Tq_{m})$, and $S_{or}(or,Tr_{h})\rightrightarrows S_{cr}(cr,Tr_{m})$.
By Definition 11, the state $S_{ou}(ou,T_{h})$ has the capacity to process $S_{oq}(oq,Tq_{h})$ and produce $S_{or}(or,Tr_{h})$ as output.
Then, by Definition 10, the states $S_{ou}(ou,T_{h})$, $S_{oq}(oq,Tq_{h})$, and $S_{or}(or,Tr_{h})$ are interpretations $E_{u}$, $E_{q}$, and $E_{r}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, all satisfying Postulate 5. Moreover, $E_{u}$ has the ability to process $E_{q}$ and produce $E_{r}$ as output.
Thus, we have the following bijective enabling mappings:
$E_{u}=S_{ou}(ou,Tu_{h})\rightrightarrows S_{M}(M,T_{m})$ (4.1)
$E_{q}$ $=S_{oq}(oq,Tq_{h})\rightrightarrows S_{cq}(cq,Tq_{m})$ (4.2)
$E_{r}$ $=S_{or}(or,Tr_{h})\rightrightarrows S_{cr}(cr,Tr_{m})$ (4.3)
By Definition 12, it follows that model M is interpretable in $\mathcal{L}$ with respect to the output $S_{cr}(cr,Tr_{m})$ under the input condition $S_{cq}(cq,Tq_{m})$. Q.E.D.
Theorem 3 fully embodies the Chinese proverb: "Whoever tied the bell must untie it." Since, intuitively, the internal logic $S_{M}(M,T_{m})$ of model M originates from the ontological state $S_{ou}(ou,Tu_{h})$ that brings the model into existence, it is thus natural to interpret the model’s output under specific input conditions using $E_{u}$, $E_{q}$, and $E_{r}$. Moreover, unlike the internal processing of $S_{M}(M,T_{m})$, which is encapsulated within the model and often opaque, the interpretations $E_{u}$, $E_{q}$, and $E_{r}$ are explicit expressions of well-formed formulas within the formal system $\mathcal{L}$ under given valuations. These expressions are more amenable to interpretation and understanding by humans, particularly by domain experts.

### v9

@@ 3.4 Information Processing and Learning Optimization in Models
User input information is formally given by:
$I_{q}=\langle oq,T_{qh},S_{oq},M_{I},t_{2},S_{M_{I}}\rangle$ (3.16)
$S_{oq}(oq,T_{qh})=\{\varphi_{q}|\varphi_{q}\text{ is a WFF about }oq\text{ and }T_{qh}\}$ (3.17)
$S_{M_{I}}(M_{I},t_{2})=\{\psi_{q}|\psi_{q}\text{ is a WFF about }M_{I}\text{ and }t_{2}\}$ (3.18)
Here, the ontology $oq$ represents the user’s intention or device, with $T_{qh}$ denoting the time at which the ontological state is formed. The state $S_{oq}(oq,T_{qh})$ is the formal representation of the user’s needs. The carrier state $S_{M_{I}}(M_{I},t_{2})$ is the physical realization, via enabling mapping, of $S_{oq}(oq,T_{qh})$ onto the model’s interface device $M_{I}$.
3.5 Information Processing, Output, and Online Learning of the Model
When the model M processes input and produces output, the information it carries can be represented as:
$I_{v}=\langle M\cup M_{I}\cup M_{O},t_{2}\cup t_{3},S_{M\cup M_{I}\cup M_{O}},M,t_{3},S_{M}\rangle$ (3.19)
Where:
$S_{M\cup M_{I}\cup M_{O}}(M\cup M_{I}\cup M_{O},t_{2}\cup t_{3})=\{\varphi_{v_{1}},\varphi_{v_{2}},\varphi_{v_{3}}\}$ (3.20)
$\varphi_{v_{1}}=State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))$ (3.21)
indicates that $\Phi^{2}(M,t_{2})$ has been assigned to M at time $t_{2}$.
$\varphi_{v_{2}}=$ $Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land$
$Output^{3}(M,t_{3},S_{M_{O}}(M_{O},t_{3})))$ (3.22)
indicates that when the input state $S_{M_{I}}(M_{I},t_{2})$ is received at $t_{2}$, M produces the output state $S_{M_{O}}(M_{O},t_{3})$ at $t_{3}$.
$\varphi_{v_{3}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
$Learnable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
$(Eq^{3}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land State^{3}(M,t_{3},\Phi^{2}(M,t_{3})))$ (3.23)
expresses that if the input information can be learned by M, then M can achieve online learning while processing and outputting information.
The physical realization of this ontological state through an enabling mapping is:
$S_{M}(M,t_{3})=\{\psi_{v_{1}},\psi_{v_{2}},\psi_{v_{3}}\}$ (3.24)
The output information is given by:
$I_{r}=\langle M\cup M_{I},t_{2},S_{M\cup M_{I}},M_{O},t_{3},S_{M_{O}}\rangle$ (3.25)
where the ontological state is:
$S_{M\cup M_{I}}(M\cup M_{I},t_{2})=\{\varphi_{r_{1}}\}$ (3.26)
with
$\varphi_{r_{1}}=Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))$ (3.27)
The formal expression of the physical state of the model’s output device $M_{O}$ at time $t_{3}$ is:
$S_{M_{O}}(M_{O},t_{3})=\{\psi_{r_{1}}\}$ (3.28)
The model can repeat the processes described in Sections 3.2 to 3.5 , iterating through cycles of training and information processing to meet user demands and continuously improve its performance.
3.6 Multi-Stage Information Causal Chains
From a systems engineering perspective, the realization of information in machine learning involves more than just five information mappings. In practice, stages such as requirements analysis, design and development, and system integration are also necessary. Specifically, in the four stages of model construction, learning and training, problem input, and answer output, and according to the corollary on information chains in Section 2.5, there is a causal and progressive process that moves from conceptual ideas to mind maps, program code, and finally to physical implementation. In this process, the earlier content serves as the ontological state, while the subsequent content becomes the carrier state, sequentially forming conceptual information, design information, and implementation information (see Figure.1).
Figure 1: Machine Learning Causality Chains with Multiple Information Mappings
As shown in Figure 1, each of these four stages encompasses four layers: conceptual ideas, mind maps, program code, and physical models. Typically, content within the same layer shares the same semantic logic and can thus be uniformly represented by a formal system. Furthermore, due to the transmissive nature of information, the innermost physical model can also be regarded as the result of enabling mappings from outer conceptual ideas or mind maps. This illustrates that the entire system comprises multiple stages, with each stage involving various combinations of information causal chains.
This framework, based on the realization of information and finite automaton states, provides a formalized representation and a comprehensive, causally progressive structure for each phase in the machine learning process. It transcends specific algorithms or models, offering a unified and universal framework, MLT-MF. In the appendix, we also present an example of its application to feedforward neural networks, which validates the correctness and applicability of the framework and supports the verification of the related theorems.
@@ 4 Several Important Theorems and Proofs
The MLT-MF framework is applicable to a range of key issues in the field of artificial intelligence, offering many unique and practical theoretical principles and methodological guidelines.
@@ 4.1 Theorem of Model Interpretability
Interpretability of complex models, such as those in deep learning, has long been one of the key challenges in artificial intelligence. However, the academic community has yet to establish a rigorous mathematical formulation of model interpretability Lipton2018 ; Doshi2017 ; Rubin2019 ; Miller2019 . In fact, interpretability should be understood as a concept that depends on the domain of discourse and the corresponding logical system. Furthermore, it should also be considered in relation to the specific inputs and outputs of the model. Based on MLT-MF, we define the following concepts:
@@
Definition 9 (Model Interpretability) Let the model M, as a finite automaton, operate over an ordered set of times $T=\{t_{i}|i=1,\ldots,n\}$, at which it undergoes assignment, input, output, and state transitions. The state of M at time $t_{i}$ is denoted as $S_{M}(M,t_{i})$, and the state of its input device $M_{I}$ at time $t_{i}$ is $S_{M_{I}}(M_{I},t_{i})$, while the state of the output device $M_{O}$ at time $t_{i+1}$ is $S_{M_{O}}(M_{O},t_{i+1})$. If there exists a formal system $\mathcal{L}$ along with three interpretation relations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$ respectively, such that
$E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$ (4.1)
@@
making $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ the respective enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, then M is said to be interpretable in $\mathcal{L}$ with respect to output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$. If for all $i=1,\cdots,n-1$, the output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in $\mathcal{L}$, then M is called interpretable in $\mathcal{L}$.
Here, $\text{Cn}(E_{M}\cup E_{M_{I}})$ denotes the logical closure of $E_{M}\cup E_{M_{I}}$. Equation 4.1 states that in the formal system $\mathcal{L}$, all formulas in $E_{M_{O}}$ are logical consequences of $E_{M}$ and $E_{M_{I}}$. Moreover, since $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are respectively enabled by $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, the formal system $\mathcal{L}$ and its three interpretations fully capture the logical and causal relationship whereby M produces output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$. This, in essence, offers a concise understanding of model interpretability.
Theorem 4 (Model Interpretability and Information Existence) Let the model M be viewed as a finite automaton operating over an ordered set of times $T=\{t_{i}|i=1,\ldots,n\}$, during which it undergoes assignment, input, output, and state transitions.
(1) For any $1\leq i\leq n-1$, a necessary and sufficient condition for the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M to be interpretable in the formal system $\mathcal{L}$ under the input $S_{M_{I}}(M_{I},t_{i})$ is that there exist $Iu(t_{i})=\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$ and $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$ representing the information carried by M and its input device $M_{I}$ at time $t_{i}$ respectively, and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$ representing the information carried by the output device $M_{O}$ at time $t_{i+1}$. The ontological states of these three pieces of information must each correspond to an interpretation in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$ must hold.
(2) A necessary and sufficient condition for M to be interpretable in $\mathcal{L}$ is that, for all $i=1,\cdots,n-1$, there exist information objects $Iu(t_{i})$, $Iq(t_{i})$, and $Ir(t_{i+1})$ whose ontological states are all interpretations in $\mathcal{L}$, and such that the ontological state of $Ir(t_{i+1})$ is contained in the logical closure of the union of the ontological states of $Iu(t_{i})$ and $Iq(t_{i})$.
Proof: For any $1\leq i\leq n-1$, suppose that the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M under the input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in the formal system $\mathcal{L}$. According to Definition 9, there exist three interpretations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ in the formal system $\mathcal{L}$, respectively over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$, such that $E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$, and $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are each the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
Let $S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))=E_{M}$, $S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))=E_{M_{I}}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))=E_{M_{O}}$. Then, according to Definition 6,we find there exist information objects $Iu(t_{i})=$ $\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$, similarly , then $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$, and $Ir(t_{i+1})=\langle or(t_{i+1}),$ $Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$, such that $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$. Thus, the necessity of the first part of the theorem is established.
On the other hand, for any $1\leq i\leq n-1$, suppose there exist some information objects $Iu(t_{i})=\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$, $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$, and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$, such that the ontological states $S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))$, $S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))$ are interpretations in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$. Then, according to Theorem 1, $E_{M}=S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))$, $E_{M_{I}}=S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))$, and $E_{M_{O}}=S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))$ are interpretations of $\mathcal{L}$ over the three finite domains $ou(t_{i})\times Tu_{h}(t_{i})$, $oq(t_{i})\times Tq_{h}(t_{i})$, and $or(t_{i+1})\times Tr_{h}(t_{i+1})$, respectively, and $E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$. At the same time, $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively. By Definition 9, the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M under the input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in the formal system $\mathcal{L}$. Therefore, the sufficiency of the first part of the theorem is established.
For all $i=1,\ldots,n-1$ , by combining Definition 9 with the first part of the theorem, the second part of the theorem follows directly. This completes the proof.